"solve the exponential equation: (3.51)^-2x/15=8
i keep getting an undefined answer. thanks for the help

Answer by Cromlix(4381)

Hi there,
(3.51)^-2x /15 = 8
(3.51)^-2x = 8 x 15
(3.51)^-2x = 120
-2x log(3.51) = log(120)
-2x = log(120)/log3.51)
-2x = 3.8128
x = 3.8128/-2
x = -1.906
Proof:
{(3.51) ^(-2 x -1.906)} /15 = 8
119.87/15 = 8
8 = 8
Hope this helps:-)